### v1

# Data Whisperer: Efficient Data Selection for Task-Specific LLM Fine-Tuning via Few-Shot In-Context Learning

Abstract
Fine-tuning large language models (LLMs) on task-specific data is essential for their effective deployment. As dataset sizes grow, efficiently selecting optimal subsets for training becomes crucial to balancing performance and computational costs. Traditional data selection methods often require fine-tuning a scoring model on the target dataset, which is time-consuming and resource-intensive, or rely on heuristics that fail to fully leverage the model’s predictive capabilities. To address these challenges, we propose Data Whisperer, an efficient, training-free, attention-based method that leverages few-shot in-context learning with the model to be fine-tuned. Comprehensive evaluations were conducted on both raw and synthetic datasets across diverse tasks and models. Notably, Data Whisperer achieves superior performance compared to the full GSM8K dataset on the Llama-3-8B-Instruct model, using just 10% of the data, and outperforms existing methods with a 3.1-point improvement and a 7.4$\times$ speedup. The code is available at https://github.com/gszfwsb/Data-Whisperer.
Data Whisperer: Efficient Data Selection for Task-Specific LLM Fine-Tuning via Few-Shot In-Context Learning
Shaobo Wang1,2thanks: Project Head (shaobowang1009@sjtu.edu.cn).  Xiangqi Jin2†  Ziming Wang2,3thanks: Xiangqi Jin and Ziming Wang are equally contributed as the second author.  Jize Wang1  Jiajun Zhang2 Kaixin Li4Zichen Wen2Zhong Li5Conghui He6Xuming Hu7Linfeng Zhang1,2thanks: Corresponding Author (zhanglinfeng@sjtu.edu.cn). 1Shanghai Jiao Tong University  2EPIC Lab, SJTU  3Nanyang Technological University 4National University of Singapore  5Microsoft Research Asia  6Shanghai AI Laboratory 7Hong Kong University of Science and Technology (Guangzhou)
1 Introduction
Fine-tuning on task-specific data has become a standard approach for adapting large language models (LLMs) to specialized tasks (Ouyang et al., 2022). For instance, continued pre-training (Gururangan et al., 2020; Chang et al., 2024) involves extending the original pre-training phase on datasets closely aligned with the target domain; instruction tuning (Zhang et al., 2023; Üstün et al., 2024) focuses on leveraging instruction-response pairs to improve the model’s ability to follow instructions; and task-specific fine-tuning enhances model performance on a particular task by using domain-specific data (Yang et al., 2024b; Zhang et al., 2024). As datasets continue to expand in size, however, a critical challenge arises: how to efficiently select the optimal training examples—a process known as data selection—to strike a balance between computational cost and model performance (Zheng et al., 2023; Paul et al., 2021; Li et al., 2023b; Zhang et al., 2024; Li et al., 2023a).
Figure 1: Comparison of the total time and performance across various data selection methods. For each method, we assessed Llama-3-8B-Instruct’s performance and time cost when utilizing 1%, 5%, and 10% of the BioInstruct dataset selected by that method.
Table 1: Comparison of Selection-to-Tuning Ratio (STR) and performance across different selection methods on various datasets using 1%, 5%, and 10% of total data points. Lower STR values indicate higher efficiency. Evaluations were performed using the Llama-3-8B-Instruct model on eight NVIDIA A100 GPUs. $\uparrow$/$\downarrow$ indicates improvement/degradation compared to random selection. “Speedup” represents the acceleration of Data Whisperer over the Nuggets (Li et al., 2023b) method. Results for Qwen and Mistral models are provided in Table 6.
Method GSM8K DialogSum BioInstruct
STR Performance STR Performance STR Performance
1% 5% 10% Avg. Score 1% 5% 10% Avg. Score 1% 5% 10% Avg. Score
Random - - - 63.77 - - - 37.66 - - - 33.89
GraNd 1.04 1.05 1.08 64.98$\uparrow$1.21 1.07 1.10 1.11 36.08$\downarrow$1.58 1.02 1.03 1.05 34.03$\uparrow$0.14
EL2N 1.05 1.07 1.12 64.44$\uparrow$0.67 1.08 1.09 1.13 32.30$\downarrow$5.36 1.02 1.03 1.06 32.14$\downarrow$1.75
CCS 1.01 1.02 1.05 64.35$\uparrow$0.58 1.01 1.03 1.05 30.65$\downarrow$7.01 1.00 1.02 1.04 34.19$\uparrow$0.30
Nuggets 1.22 1.23 1.26 66.07$\uparrow$2.30 2.42 2.46 2.53 36.25$\downarrow$1.41 0.46 0.47 0.49 32.51$\downarrow$1.38
Data Whisperer 0.13 0.14 0.17 68.14$\uparrow$4.37 0.12 0.15 0.25 40.40$\uparrow$2.74 0.03 0.04 0.06 36.46$\uparrow$2.57
Speedup 9.38$\times$ 8.79$\times$ 7.41$\times$ - 20.17$\times$ 16.40$\times$ 10.12$\times$ - 15.33$\times$ 11.75$\times$ 8.17$\times$ -
In this paper, we focus on data selection for task-specific datasets, aiming to extract the most informative subset from the original data to achieve nearly lossless performance compared to using the entire dataset for fine-tuning [^1] We address scenarios where no more advanced LLMs are available beyond the model that is to be fine-tuned, e.g., GPT-4 (OpenAI et al., 2023).. Although current data selection methods have demonstrated remarkable performance, they often suffer from significant inefficiencies. For instance, as shown in Figure 1, previous state-of-the-art (SOTA) methods can take more time than directly using the entire BioInstruct dataset to fine-tune a Llama-3-8B-Instruct model, even when working with just 1% of the original dataset. This inefficiency arises because these approaches depend on a fine-tuned model on the target dataset for scoring (Li et al., 2023a; Xia et al., 2024), typically requiring the same architecture as the model to be fine-tuned.
To this end, we first critically reevaluate the effectiveness of existing selection methods. To quantitatively and fairly assess the effectiveness of each method, we introduce the Selection-to-Tuning Ratio (STR), which is defined as the ratio of time spent on selection to the time required for fine-tuning the model on the entire dataset. Formally, let $t_{p}(\tau,\rho)$ represent the time associated with a selection method $\tau$ with a budget subset ratio $\rho$, and let $t_{ft}$ denote the corresponding fine-tuning time for the entire dataset. The STR is given by:
$\text{STR}(\tau)=\frac{t_{p}(\tau,\rho)}{t_{ft}}.$ (1)
Intuitively, for a data selection method $\tau$ to be considered practically efficient, its STR should ideally be less than 1. This condition ensures that the time spent on selection aligns with the benefits obtained during fine-tuning. However, as shown in Table 1, existing methods often exhibit an STR greater than 1, which constitutes a significant bottleneck in the data selection process, thereby severely limiting the scalability of these methods for large datasets and models.
Figure 2: Illustration of the proposed Data Whisperer. $n_{d}$ and $n_{q}$ denote the number of demonstrations and queries in ICL. The pipeline consists of two main steps: (I) Few-shot In-Context Learning. A set of demonstration and query examples is randomly sampled from the initial dataset, and an ICL prompt is constructed with a fixed instruction. The LLM to be fine-tuned generates answers for all query examples, and the average evaluation score is computed using the ground truth answers. (II) Context-Aware Weighting. During each iteration of few-shot ICL, we weight the scores of the demonstration examples based on their attention scores, which quantify their influence on the queries. The scores are updated until all samples are scored. The final score for each sample is the average score across its appearances in multiple demonstration sets. Finally, we select the top-k samples from the dataset.
To address the inefficiencies of existing data selection methods, we introduce Data Whisperer, a training-free, attention-based, and context-aware approach. Traditional methods typically rely on an additional LLM that is fine-tuned on the target dataset to provide scoring for data selection. In contrast, our approach takes advantage of the model’s inherent predictive capabilities, inspired by recent theories that equate in-context learning (ICL) with fine-tuning in LLMs (Dai et al., 2022; Mosbach et al., 2023). As shown in Figure 2, we score each training sample using few-shot ICL, where the sample itself serves as a demonstration example. The model’s performance, measured by its ability to respond to associated queries based on these demonstration examples, yields a raw score for each sample. To improve efficiency, this scoring is performed with a weak-to-strong strategy, i.e., using a less powerful pre-trained model from the same model family.
In typical ICL, context-awareness (e.g., the order of examples within the context window) can influence the model’s performance due to its reliance on sequence-based attention mechanisms (Guo et al., 2024; Bhope et al., 2025). To mitigate this inherent order sensitivity, we weight the demonstration scores according to their respective attention scores, as shown in Figure 2. This context-aware weighting refines the selection process by ensuring that the final score for each training sample not only reflects its raw performance but also accounts for its contextual significance within the task. The final score for each sample is computed as the average of these weighted scores, which provides a more accurate and nuanced evaluation of each data point’s contribution to the model’s learning process.
Our contributions can be summarized as follows:
  1. 1.
    We critically examine existing data selection approaches and introduce the Selection-to-Tuning Ratio, a novel metric that quantifies the efficiency of these methods. We observed that all prior data selection methods are more inefficient than fine-tuning the LLM with the entire dataset.
  2. 2.
    We propose Data Whisperer, an effective, training-free and attention-based method. Unlike previous approaches, our method eliminates the need to fine-tune a separate scoring model on the target dataset, ensuring greater efficiency.
  3. 3.
    Data Whisperer integrates seamlessly with weak-to-strong few-shot ICL schemes, enabling effective performance even when a weaker model within the same model family is employed for ICL. This enhances both the scalability and efficiency of our method.
  4. 4.
    Comprehensive experiments are conducted on both real and synthetic datasets across various selection budget ratios, including BioInstruct (Tran et al., 2024), DialogSum (Chen et al., 2021), and GSM8K (Cobbe et al., 2021). We observe that Data Whisperer consistently outperforms previous SOTA methods, particularly in smaller data scenarios, while achieving faster selection times.
2 Related Work
2.1 In-Context Learning
In-Context Learning (ICL) is a powerful task adaptation technique that does not require modifying the weights of a pre-trained model (Brown et al., 2020; Olsson et al., 2022; Laskin et al., 2022). Instead, ICL adapts the model to new tasks by conditioning it on a sequence of demonstration pairs, where each demonstration consists of input-output pairs formatted according to a predefined template. This process guides the model in understanding the task. During inference, ICL involves presenting the model with a series of demonstrations followed by a query set, with the model expected to predict the corresponding labels for the query data points based on the context.
Recent studies have explored the theoretical links between ICL and gradient descent, suggesting an implicit relationship between ICL and fine-tuning (Dai et al., 2022; Mosbach et al., 2023; Deutch et al., 2023; Zhou et al., 2024; Chen et al., 2024). One approach, Nuggets (Li et al., 2023b), uses one-shot learning for selection in instruction-tuning tasks with log probability scores, but is limited by the computational inefficiency of one-shot learning. In contrast, our work employs attention-based few-shot ICL on the pre-trained model, before any fine-tuning, to directly measure performance scores for task-specific fine-tuning. This approach leverages the model’s inherent knowledge for data selection, offering a more efficient and scalable solution by linking attention-aware ICL predictions to the fine-tuning process.
2.2 Data Selection
Data selection is critical in deep learning, aiming to identify high-quality data that enhances training efficiency while maintaining comparable (or nearly lossless) performance. Traditional methods rely on heuristic metrics, with two primary strategies: (I) importance-based selection, which targets challenging or essential samples (Raffel et al., 2020; Paul et al., 2021; Mirzasoleiman et al., 2020; Killamsetty et al., 2021; Tan et al., 2024), and (II) similarity- or diversity-based selection, which selects samples based on their similarity to others or their representativeness of the feature space (Bukharin and Zhao, 2023; Xia et al., 2024; Zheng et al., 2023; Wei et al., 2021; Yu et al., 2024).
For LLMs, these methods face two challenges: (I) the unreliability of heuristic metrics in high-dimensional spaces, and (II) the computational cost of fine-tuning additional scoring models. In contrast, our approach directly integrates data selection with the pre-trained model’s predictions using ICL, eliminating the need for fine-tuning or heuristics.
3 Method: Data Whisperer
3.1 Preliminary: Data Selection in LLM Task-Specific Fine-Tuning
Given a large and diverse task dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{|\mathcal{D}|}$, the objective of data selection is to identify a representative subset $\mathcal{D^{\prime}}\subseteq\mathcal{D}$, with a budget ratio $\rho=|\mathcal{D^{\prime}}|/|\mathcal{D}|<1$, for fine-tuning a pre-trained language model $\mathcal{M}_{p}$. The fine-tuned model, $\mathcal{M}_{f}$, is then evaluated on the test set $\mathcal{D}_{\textrm{test}}$. The selected subset $\mathcal{D^{\prime}}$ should ensure that the performance of the fine-tuned model $\mathcal{M}_{f}(\mathcal{D}^{\prime})$ remains comparable to that of a model fine-tuned on the entire dataset $\mathcal{M}_{f}(\mathcal{D})$ when evaluated on $\mathcal{D}_{\textrm{test}}$. The data selection objective is formulated as:
$\min_{\mathcal{D^{\prime}}\subseteq\mathcal{D},|\mathcal{D^{\prime}}|=\rho\cdot|\mathcal{D}|}\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{test}}}\ell(x,y;\mathcal{M}_{f}(\mathcal{D}^{\prime})),$ (2)
where $\ell(\cdot)$ is the task-specific loss function (e.g., cross-entropy). The key challenge in effective coreset selection is to efficiently identify the coreset $\mathcal{D}^{\prime}$ for different selection ratios $\rho$.
3.2 ICL as a Data Selector for LLM Fine-Tuning
In contrast to conventional methods that rely on handcrafted heuristics of a fine-tuned model $\mathcal{M}_{f}$, our approach, Data Whisperer, utilizes the intrinsic ICL capabilities of the pre-trained model $\mathcal{M}_{p}$ for data selection. At each iteration, we randomly sample $n_{d}$ demonstration examples $\mathcal{D}_{d}=\{(x_{d}^{(1)},y_{d}^{(1)}),\dots,(x_{d}^{(n_{d})},y_{d}^{(n_{d})})\}$, and $n_{q}$ query examples $\mathcal{D}_{q}=\{x_{q}^{(1)},\dots,x_{q}^{(n_{q})}\}$, where $\mathcal{D}_{d}\cap\mathcal{D}_{q}=\emptyset$. Combined with a fixed instruction $I$, the total context $C$ is obtained, which is defined as:
$C=\{I,(x_{d}^{(1)},y_{d}^{(1)}),\dots,(x_{d}^{(n_{d})},y_{d}^{(n_{d})})\},$ (3)
The whole ICL process is formalized as:
$\hat{y}_{q}^{(1)},\dots,\hat{y}_{q}^{(n_{q})}=\mathcal{M}_{p}(C,x_{q}^{(1)},\dots,x_{q}^{(n_{q})}).$ (4)
Using performance metrics $f$ (e.g., average accuracy, or ROUGE-L (Lin, 2004), etc.), we compare the predicted outputs $\hat{y}_{q}^{(1)},\dots,\hat{y}_{q}^{(n_{q})}$ with the ground truth labels $y_{q}^{(1)},\dots,y_{q}^{(n_{q})}$ to compute the average performance score $s$:
$s=\frac{1}{n_{q}}\sum_{j=1}^{n_{q}}f(\hat{y}_{q}^{(j)},y_{q}^{(j)}).$ (5)
This score is assigned to each sample in the demonstration set, and the process is repeated across multiple iterations. The final score for each sample is the average of its scores from all repetitions.
3.3 Context-Aware Weighting
To mitigate potential order sensitivity in few-shot ICL, we introduce a context-aware weighting mechanism for scoring the demonstration set, as shown in Figure 2. This mechanism leverages the self-attention scores from a specific layer $l$ of the pre-trained model $\mathcal{M}_{p}$, across all attention heads $h$.
We focus on the attention scores corresponding to the first prediction token of a fixed layer $l$. Let $A^{(h)}$ denote the self-attention matrix of layer $l$. For each demonstration example $(x_{d}^{(i)},y_{d}^{(i)})$, we extract the submatrix $A^{(h)}_{(x_{d}^{(i)},y_{d}^{(i)})}$ from $A^{(h)}$, which contains the attention scores between this demonstration and all query examples $x_{q}^{(1)},\dots,x_{q}^{(n_{q})}$. The sum of these attention scores quantifies the interaction between the demonstration and the query examples. The weight for each demonstration example, across all attention heads, is computed as:
$w_{(x_{d}^{(i)},y_{d}^{(i)})}=\sum_{h}\mathbf{1}^{\top}A_{(x_{d}^{(i)},y_{d}^{(i)})}^{(h)}\mathbf{1},$ (6)
where $A_{(x_{d}^{(i)},y_{d}^{(i)})}^{(h)}$ is the attention submatrix for head $h$, and the summation is performed across all heads. To account for variations in the length of demonstration examples, we normalize the summed attention scores by the length of each demonstration.
This process is repeated for each demonstration in every iteration until all samples are scored, yielding the scoring set $\mathcal{S}$. If any demonstration example is scored more than once, the scores are averaged. Finally, we select the Top-k samples as follows:
$\mathcal{D}^{\prime}=\text{Top-k}(\mathcal{D};\mathcal{S}),$ (7)
where $k=\lfloor\rho\cdot|\mathcal{D}|\rfloor$ denotes the coreset size. The detailed algorithm of Data Whisperer is demonstrated in Algorithm 1.
3.4 Theoretical Analysis
We now provide a theoretical analysis of Data Whisperer and demonstrate that our approach effectively performs implicit data selection for task-specific fine-tuning.
Analysis of ICL’s Query Prediction. In ICL, the model adjusts its predictions based on the attention assigned to a set of demonstration examples. Let $x\in\mathbb{R}^{d_{\text{in}}}$ represent the input of a query token $t$, and $q=W_{Q}x$ denote the attention query vector. The attention result for a specific head is formulated as:
$\mathcal{M}_{p}(q)=\textrm{Attn}(V,K,q)$ (8)
$=W_{V}[X_{d};X_{q,<t}]\sigma\left(\frac{(W_{K}[X_{d};X_{q,<t}]^{\top}q)}{\sqrt{d}}\right),$
where $\sigma$ denotes the softmax operator, and $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ are the projection matrices for the attention queries, keys, and values, and $\sqrt{d}$ is the scaling factor. Here, $X_{q,<t}$ represents the input representations of the query tokens before token $t$, $X_{d}$ represents the input representations of the demonstration tokens, and $[X_{d};X_{q,<t}]$ represents the concatenation of these matrices. For qualitative analysis, we approximate the standard attention mechanism by relaxing it to linear attention, removing both the softmax operation and the scaling factor:
$\mathcal{M}_{p}(q)\approx W_{V}[X_{d};X_{q,<t}](W_{K}[X_{d};X_{q,<t}])^{\top}q$ (9)
$=\underbrace{W_{V}X_{q,<t}(W_{K}X_{q,<t})^{\top}q}_{\text{zero shot input}}+\underbrace{W_{V}X_{d}(W_{K}X_{d})^{\top}q}_{\text{ICL examples input}}$
$=W_{\textrm{zsl}}q+\sum_{i}(W_{V}x_{d}^{(i)})\left((W_{K}x_{d}^{(i)})^{\top}q\right)$
$=W_{\textrm{zsl}}q+\sum_{i}\left((W_{V}x_{d}^{(i)})\otimes(W_{K}x_{d}^{(i)})\right)q$
$=(W_{\textrm{zsl}}+\Delta W_{\textrm{icl}})q$
In the above derivation, we observe that the attention to the demonstration tokens, $X_{d}$, can be viewed as an implicit parameter update $\Delta W_{\text{icl}}$, which modifies the zero-shot weights $W_{\text{zsl}}$ to adapt to the task at hand. This update is essentially driven by the in-context learning examples and their relevance to the query tokens.
Analysis of Fine-Tuning for Query Prediction. In fine-tuning, model explicitly updates its key and value projections, $W_{K}$ and $W_{V}$, through backpropagation to improve task performance. After fine-tuning, the model’s attention can be expressed as:
$\mathcal{M}_{f}(q)$ $=(W_{V}+\Delta W_{V})X_{q,<t}$ (10)
$\cdot X_{q,<t}^{\top}(W_{K}+\Delta W_{K})^{\top}q$
$=\underbrace{W_{V}X_{q,<t}(W_{K}X_{q,<t})^{\top}q}_{\text{zero shot input}}$
$+\Delta W_{V}X_{q,<t}(W_{K}X_{q,<t})^{\top}q$
$+\Delta W_{V}X_{q,<t}(\Delta W_{K}X_{q,<t})^{\top}q$
$+\Delta W_{V}X_{q,<t}(\Delta W_{K}X_{q,<t})^{\top}q$
$=(W_{\textrm{zsl}}+\Delta W_{\textrm{ft}})q.$
Here, $\Delta W_{K}$ and $\Delta W_{V}$ represent the parameter updates to the key and value projections $W_{K}$ and $W_{V}$, respectively, which are learned via backpropagation from task-specific training objectives. The update $\Delta W_{\text{ft}}$ corresponds to the changes in $W_{\text{zsl}}$ introduced by the fine-tuning process.
Connecting ICL Data Selection and Fine-Tuning. Based on the previous analysis, both ICL and FT follow a similar structure for task adaptation. Both methods modify the attention parameters $W_{K}$ and $W_{V}$, but the difference lies in how the model updates them. In ICL, as shown in Eq. (9), the demonstration examples $x_{d}^{(i)}$ impact the query prediction by adjusting the attention weights. These weights, $(W_{V}x_{d}^{(i)})\otimes(W_{K}x_{d}^{(i)})$, determine relevance of each example. In contrast, as shown in Eq. (10), FT explicitly refines $W_{K}$ and $W_{V}$ through gradients from the task-specific loss function.
Given the structural similarity between ICL and FT, we infer that using ICL for data selection is valid for FT. By selecting and weighting the most relevant examples in ICL, we achieve the same performance gains as fine-tuning, but without the need for explicit updates. Thus, Data Whisperer efficiently performs data selection through ICL, identifying the most relevant data points and adapting the model with minimal computational overhead.
Table 2: Evaluation results of different data selection methods on the GSM8k, DialogSum, and BioInstruct datasets. The model was fine-tuned on a coreset comprising 1%, 5%, and 10% of the original dataset size. $\uparrow$ indicates an improvement over random selection, while $\downarrow$ indicates a degradation compared to random selection.
Method GSM8k DialogSum BioInstruct
1% 5% 10% 1% 5% 10% 1% 5% 10%
Llama-3-8B-Instruct (Zero-shot) 57.49 18.57 13.80
+ Random 56.95 64.71 69.66 31.66 39.88 41.45 28.38 34.61 38.70
+ GraNd (Paul et al., 2021) 57.23$\uparrow$0.28 67.65$\uparrow$2.94 70.05$\uparrow$0.39 26.02$\downarrow$5.64 40.32$\uparrow$0.44 41.90$\uparrow$0.45 28.73$\uparrow$0.35 36.47$\uparrow$1.86 36.88$\downarrow$1.82
+ EL2N (Paul et al., 2021) 60.16$\uparrow$3.21 61.50$\downarrow$3.21 71.66$\uparrow$2.00 23.05$\downarrow$8.61 35.14$\downarrow$4.74 38.71$\downarrow$2.74 28.29$\downarrow$0.09 31.52$\downarrow$3.09 36.62$\downarrow$2.08
+ CCS (Zheng et al., 2023) 60.70$\uparrow$3.75 63.64$\downarrow$0.07 68.72$\downarrow$0.94 23.48$\downarrow$8.18 27.34$\downarrow$12.54 41.12$\downarrow$0.33 29.25$\uparrow$0.87 36.71$\uparrow$2.10 36.61$\downarrow$2.09
+ Nuggets (Li et al., 2023b) 60.17$\uparrow$3.22 68.65$\uparrow$3.94 69.39$\downarrow$0.27 31.77$\uparrow$0.11 36.85$\downarrow$3.03 40.13$\downarrow$1.32 22.95$\downarrow$5.43 36.41$\uparrow$1.80 38.18$\downarrow$0.52
+ Data Whisperer (ours) 62.57$\uparrow$5.62 69.65$\uparrow$4.94 72.46$\uparrow$2.80 38.05$\uparrow$6.39 40.96$\uparrow$1.08 42.18$\uparrow$0.73 32.29$\uparrow$3.91 37.27$\uparrow$2.66 39.20$\uparrow$0.50
Whole Dataset 71.39 43.33 40.21
Qwen-2.5-7B-Instruct (Zero-shot) 21.23 22.10 17.30
+ Random 21.51 79.28 80.08 27.31 36.68 39.99 30.65 36.62 38.30
+ GraNd (Paul et al., 2021) 21.98$\uparrow$0.47 82.09$\uparrow$2.81 83.76$\uparrow$3.68 30.61$\uparrow$3.30 37.77$\uparrow$1.09 40.50$\uparrow$0.51 23.70$\downarrow$6.95 35.89$\downarrow$0.73 38.27$\downarrow$0.03
+ EL2N (Paul et al., 2021) 22.13$\uparrow$0.62 77.41$\downarrow$1.87 82.62$\uparrow$2.54 32.57$\uparrow$5.26 38.90$\uparrow$2.22 39.45$\downarrow$0.54 30.36$\downarrow$0.29 36.30$\downarrow$0.32 37.26$\downarrow$1.04
+ CCS (Zheng et al., 2023) 21.33$\downarrow$0.18 80.75$\uparrow$1.47 81.95$\uparrow$1.87 31.69$\uparrow$4.38 36.15$\downarrow$0.53 38.05$\downarrow$1.94 23.73$\downarrow$6.92 34.53$\downarrow$2.09 37.34$\downarrow$0.96
+ Nuggets (Li et al., 2023b) 22.70$\uparrow$1.19 82.09$\uparrow$2.81 83.42$\uparrow$3.34 32.11$\uparrow$4.80 37.54$\uparrow$0.86 40.23$\uparrow$0.24 30.25$\downarrow$0.40 36.33$\downarrow$0.29 38.28$\downarrow$0.02
+ STAFF (Zhang et al., 2024) 22.52$\uparrow$1.01 82.22$\uparrow$2.94 83.49$\uparrow$3.41 30.16$\uparrow$2.85 40.84$\uparrow$4.16 40.54$\uparrow$0.55 23.64$\downarrow$7.01 37.08$\uparrow$0.46 38.20$\downarrow$0.10
+ Data Whisperer (ours) 24.45$\uparrow$2.94 83.16$\uparrow$3.88 85.03$\uparrow$4.95 32.95$\uparrow$5.64 40.95$\uparrow$4.27 43.00$\uparrow$3.01 34.93$\uparrow$4.28 37.57$\uparrow$0.95 38.85$\uparrow$0.55
Whole Dataset 85.43 43.79 40.71
Mistral-Nemo-Instruct (Zero-shot) 29.41 19.39 13.12
+ Random 32.35 54.28 64.44 19.48 37.37 40.98 13.25 24.98 36.76
+ GraNd (Paul et al., 2021) 31.55$\downarrow$0.80 56.02$\uparrow$1.74 67.38$\uparrow$2.94 19.86$\uparrow$0.38 36.52$\downarrow$0.85 41.79$\uparrow$0.81 17.24$\uparrow$3.99 25.76$\uparrow$0.78 30.05$\downarrow$6.71
+ EL2N (Paul et al., 2021) 30.43$\downarrow$1.92 57.65$\uparrow$3.37 67.14$\uparrow$2.70 19.97$\uparrow$0.49 35.88$\downarrow$1.49 36.40$\downarrow$4.58 14.93$\uparrow$1.68 26.20$\uparrow$1.22 30.15$\downarrow$6.61
+ CCS (Zheng et al., 2023) 32.35$\downarrow$0.00 53.61$\downarrow$0.67 66.18$\uparrow$1.74 18.55$\downarrow$0.93 34.57$\downarrow$2.80 42.14$\uparrow$1.16 14.42$\uparrow$1.17 28.12$\uparrow$3.14 36.55$\downarrow$0.21
+ Nuggets (Li et al., 2023b) 32.09$\downarrow$0.26 58.29$\uparrow$4.01 68.79$\uparrow$4.35 19.52$\uparrow$0.04 37.54$\uparrow$0.17 40.52$\downarrow$0.46 20.40$\uparrow$7.15 27.40$\uparrow$2.42 36.56$\downarrow$0.20
+ STAFF (Zhang et al., 2024) 31.08$\downarrow$1.27 63.18$\uparrow$8.90 67.91$\uparrow$3.47 19.49$\uparrow$0.01 36.55$\downarrow$0.82 42.08$\uparrow$1.10 13.26$\uparrow$0.01 25.40$\uparrow$0.42 36.44$\downarrow$0.32
+ Data Whisperer (ours) 32.63$\uparrow$0.28 65.91$\uparrow$11.63 74.32$\uparrow$9.88 21.52$\uparrow$2.04 41.48$\uparrow$4.11 43.36$\uparrow$2.38 21.54$\uparrow$8.29 29.98$\uparrow$5.00 38.37$\uparrow$1.61
Whole Dataset 75.00 43.69 40.24
Table 3: Evaluation results of different data selection methods on the synthetic dataset, generated from DialogSum dataset. The model was fine-tuned on a coreset comprising 5%, 10%, and 25% of the synthetic dataset size.
Method Llama-3-8B-Instruct Qwen-2.5-7B-Instruct Mistral-Nemo-Instruct-2407
5% 10% 25% 100% 5% 10% 25% 100% 5% 10% 25% 100%
Random 30.21 31.26 33.05 35.31 26.15 27.46 31.31 34.55 19.24 32.44 33.09 34.36
GraNd 27.04$\downarrow$3.17 31.38$\uparrow$0.12 33.34$\uparrow$0.29 28.00$\uparrow$1.85 29.79$\uparrow$2.33 30.59$\downarrow$0.72 20.03$\uparrow$0.79 32.28$\downarrow$0.16 33.05$\downarrow$0.04
EL2N 28.67$\downarrow$1.54 31.26$\downarrow$0.00 32.78$\downarrow$0.27 27.60$\uparrow$1.45 27.33$\downarrow$0.13 34.91$\uparrow$3.60 20.96$\uparrow$1.72 30.92$\downarrow$1.52 32.42$\downarrow$0.67
CCS 30.19$\downarrow$0.02 32.75$\uparrow$1.49 33.77$\uparrow$0.72 26.41$\uparrow$0.26 33.59$\uparrow$6.13 33.83$\uparrow$2.52 19.77$\uparrow$0.53 32.85$\uparrow$0.41 34.07$\uparrow$0.98
Nuggets 30.39$\uparrow$0.18 30.83$\downarrow$0.43 33.84$\uparrow$0.79 28.40$\uparrow$2.25 28.02$\uparrow$0.56 32.05$\uparrow$0.74 19.98$\uparrow$0.74 31.98$\downarrow$0.46 32.46$\downarrow$0.63
STAFF - - - 27.33$\uparrow$1.18 29.64$\uparrow$2.18 31.79$\uparrow$0.48 20.88$\uparrow$1.64 31.11$\downarrow$1.33 31.33$\downarrow$1.76
Data Whisperer 32.15$\uparrow$1.94 32.81 $\uparrow$1.55 34.07$\uparrow$1.02 31.27$\uparrow$5.12 34.04$\uparrow$6.58 35.20$\uparrow$3.89 22.35$\uparrow$3.11 35.08$\uparrow$2.64 35.49$\uparrow$2.40
4 Experiments
4.1 Experimental Setup
Datasets. We conducted experiments across four datasets, including three real-world and one synthetic dataset, each corresponding to a distinct downstream task. Specifically, we used the following datasets: (i) the BioInstruct dataset (Tran et al., 2024) for biomedical question answering, (ii) the DialogSum dataset (Chen et al., 2021) for dialogue summarization, (iii) GSM8K (Cobbe et al., 2021) for mathematical reasoning, and (iv) a synthetic variant of DialogSum. Please see Appendix A.2 for detailed prompt designs.
Models. For each task, we evaluated the performance of three widely-used large language models (LLMs): Llama-3-8B-Instruct (Dubey et al., 2024), Qwen-2.5-7B-Instruct (Yang et al., 2024a; Team, 2024), and Mistral-Nemo-Instruct-2407 (Jiang et al., 2023). To explore the potential of using a weaker model for ICL, Data Whisperer also incorporated two smaller LLMs: Qwen-2.5-3B-Instruct (Yang et al., 2024a; Team, 2024) and Mistral-7B-Instruct-v0.2 (Jiang et al., 2023). Notably, since Llama-3-8B-Instruct does not have a smaller variant within its model family, it was excluded from the weak-to-strong analysis.
Baselines. We compared Data Whisperer with several state-of-the-art data selection techniques: (i) Random Selection, which randomly samples subsets from the dataset, (ii) GraNd (Paul et al., 2021), which selects samples based on large gradient norms, (iii) EL2N (Paul et al., 2021), which prioritizes samples with large discrepancies between model predictions and ground truth, (iv) CCS (Zheng et al., 2023), which balances data coverage and importance, (v) Nuggets (Li et al., 2023b), which employs one-shot learning of a fine-tuned model to identify high-quality examples, and (vi) STAFF (Zhang et al., 2024), which estimates gradient effort scores with a smaller model and validates them on the target LLM to allocate budget across different model regions.
Evaluation. All experiments were performed on 8 NVIDIA A100 GPUs. For model fine-tuning, we utilized LoRA (Hu et al., 2021) and performed 3 times. Performance on the BioInstruct and DialogSum datasets was assessed using the ROUGE-L metric (Lin, 2004) during both the data selection and fine-tuning stages. For the GSM8K dataset, we computed performance based on the Exact Match (EM) metric by comparing the model’s generated answers against the ground truth.
Synthetic Data Generation. We generated a synthetic variant of the DialogSum dataset using the Llama-3-8B-Instruct model. For every five demonstration samples from the original dataset, we prompted the model to generate one synthetic sample, ensuring alignment with the original dialogue structure. The generated samples were manually reviewed and filtered to remove incoherent, factually incorrect, or misformatted data. See Appendix A.3 for details on synthetic data generation.
Table 4: Ablation study on the number of demonstrations $n_{d}$ and queries $n_{q}$ in Data Whisperer. We observed that selecting a moderate number of $n_{q}$ and $n_{d}$ often yields a better performance for Data Whisperer.
Dataset Ratio Llama-3-8B-Instruct Qwen-2.5-7B-Instruct
$n_{d}=5$ $n_{q}=3$ $10$ $5$ $15$ $10$ $5$ $3$ $10$ $5$ $15$ $10$
GSM8K 0% 57.49 21.23
1% 62.57 60.96 61.23 22.43 24.45 23.75
5% 68.18 69.65 66.56 81.95 83.01 80.95
10% 70.95 72.46 70.72 83.75 84.89 85.03
DialogSum 0% 18.57 22.10
1% 36.11 36.47 36.45 32.49 32.20 32.14
5% 38.15 40.96 38.43 38.02 40.95 38.46
10% 41.27 42.18 41.42 37.88 43.00 40.31
BioInstruct 0% 13.80 17.30
1% 32.29 32.24 30.45 32.34 34.93 32.07
5% 32.96 37.27 34.01 36.74 37.14 36.64
10% 38.98 39.20 37.67 38.20 38.85 38.16
4.2 Main Results and Key Observations
$\bullet$ Smaller Dataset, Comparable Performance with the Whole Dataset. For real datasets, as shown in Table 2, results demonstrate that Data Whisperer performs remarkably well on smaller datasets, achieving comparable performance to full dataset across various selection ratios. For instance, on the GSM8K dataset, fine-tuning with 10% of data yields even better performance to the model fine-tuned on the entire dataset. For synthetic dataset in Table 3, two out of three models outperform the models fine-tuned on the entire dataset.
$\bullet$ Same Size, Better Performance with SOTA Baseline Methods. Compared to SOTA methods, Data Whisperer demonstrates consistent superiority across varying dataset sizes. On real datasets, as illustrated in Table 2, Data Whisperer achieves higher accuracy. For instance, on 10% data of DialogSum with Qwen-2.5-7B-Instruct, Data Whisperer attains an accuracy of 43.00, surpassing the previous SOTA method, STAFF, by a significant margin of 2.46. Similarly, on synthetic datasets, as shown in Table 3, Data Whisperer consistently delivers the best performance across all evaluated models and data proportions, underscoring its robust generalization capabilities. Notably, with the Qwen-2.5-7B-Instruct model on 5% of the data, Data Whisperer achieves an accuracy of 31.27, outperforming the prior SOTA method, Nuggets, by a remarkable 2.87 points.
Table 5: Results of different attention layers for context-aware weighting, and the impact of context-aware weighting (w/ vs. w/o) on Llama-3-8B-Instruct.
Dataset Ratio Random Data Whisperer
w/o Attn Shallow Intermediate Deep
GSM8K 0% 57.49
1% 56.95 60.56 60.73 60.96 62.30
5% 64.71 68.45 68.98 69.65 68.71
10% 69.66 71.72 71.25 72.46 71.79
DialogSum 0% 18.57
1% 31.66 34.35 37.97 36.47 38.05
5% 39.88 40.29 39.96 40.96 40.08
10% 41.45 41.91 41.94 42.18 41.80
BioInstruct 0% 13.80
1% 28.38 30.51 29.27 32.24 29.44
5% 34.61 36.38 36.36 37.27 36.32
10% 38.70 38.83 38.86 39.20 38.79
4.3 Ablation Study
$\bullet$ Sensitivity of demonstration and query numbers $n_{d}$ and $n_{q}$. An interesting question is how the number of demonstration examples ($n_{d}$) and query examples ($n_{q}$) affects the performance of Data Whisperer. To explore this, we varied the values of $n_{d}$ and $n_{q}$ and observed their impact across all three datasets. As shown in Table 4, results indicate that while increasing $n_{d}$ or $n_{q}$ improves the model’s performance at first, the effect tends to plateau after a certain threshold. This suggests that there exists an optimal balance between the number of demonstration and query examples, beyond which the improvements are marginal, demonstrating the robustness of Data Whisperer. We used $n_{d}=10$ and $n_{q}=5$ by default.
Figure 3: Ablation on the weak-to-strong scalability. For Qwen models, we used Qwen-2.5-3B-Instruct to perform Data Whisperer for Qwen-2.5-7B-Instruct. For Mistral models, we used Mistral-7B-Instruct-v0.2 to perform Data Whisperer for Mistral-Nemo-Instruct-2407. Weaker models were not fine-tuned on the task dataset. We found that using a weaker model does not significantly affect performance and provides a more efficient solution (measured by STR). Best viewed in color.
$\bullet$ Ablation of Attention Layers. The effectiveness of the context-aware weighting mechanism depends on the selected attention layer for scoring. We conducted an ablation study on the Llama-3-8B-Instruct model, evaluating shallow (Layer 5), intermediate (Layer 13), and deep (Layer 30) layers. As shown in Table 5, the intermediate layer (e.g., Layer 13) generally yielded more stable results, though all layers outperformed random selection. This suggests that Data Whisperer benefits from multiple attention layers, with intermediate layers providing more relevant contextual information for task-specific fine-tuning. These findings align with previous research showing that intermediate layers are crucial for semantic interpretation, processing moderately complex concepts, and providing contextual information for tasks requiring understanding (Wendler et al., 2024; Jin et al., 2025; Li et al., 2024).
$\bullet$ Effect of weak-to-strong scoring. We also investigated the impact of weak-to-strong scoring, where a weaker model is used to select data for fine-tuning a stronger model. As shown in Figure 3, results indicate that using a weaker model does not significantly impact the overall performance of Data Whisperer, while providing a more efficient solution with a lower STR. It demonstrates that Data Whisperer is scalable across different model sizes and highlights its potential for efficient fine-tuning, even with limited computational resources.
5 Discussion: What Samples Do Task-Specific Fine-Tuning Prefer?
To study what kinds of samples task-specific fine-tuning favors, we analyzed the perplexity of both the full dataset and the selected 1% subset across three different task datasets, including BioInstruct, GSM8K, and DialogSum. Specifically, we used GPT-4o mini (OpenAI et al., 2023) to score the perplexity of both the full data and the 1% selected subset. As shown in Figure 4, our findings show that the selected 1% of data consistently exhibits lower perplexity (i.e., easier data) across all three datasets. This aligns with previous findings in (Sorscher et al., 2022), which suggest that, in small data scenarios, the model tends to prefer easier data to boost performance. Lower perplexity data tends to be more predictable and less ambiguous, allowing the model to learn task-specific patterns more efficiently when fine-tuned on such samples. This observation further supports the notion that task-specific fine-tuning benefits from focusing on simpler, high-confidence examples, especially in data-limited settings.
Figure 4: Perplexity distribution of all samples and the selected 1% samples by Data Whisperer. Perplexity scores are calculated using the GPT-4o-mini model API. The results show that Data Whisperer prefers selecting easier samples, which aligns with the theory in (Sorscher et al., 2022) that suggests selecting the easiest samples is optimal in small data scenarios (e.g., 1% data).
6 Conclusion
In this paper, we reevaluate existing data selection methods through the lens of the Selection-to-Tuning Ratio, revealing that many traditional approaches fall short in practical scenarios. To address this gap, Data Whisperer introduces an efficient, training-free, attention-based approach to data selection. By leveraging the theoretical connection between ICL and fine-tuning, Data Whisperer eliminates the need for fine-tuning an additional LLM on the target dataset for scoring. Specifically, we use few-shot ICL for data selection, with randomly sampled demonstration and query examples drawn from the initial dataset. Moreover, to mitigate the inherent order sensitivity in ICL, our method incorporates a context-aware weighting strategy based on attention scores. In conclusion, Data Whisperer provides a scalable, efficient solution for data selection in task-specific LLM fine-tuning, offering significant potential for enhancing both efficiency and performance.
7 Limitation
While Data Whisperer demonstrates strong performance across a range of tasks and datasets, there are opportunities for further exploration. The attention-based scoring mechanism relies on access to the internal attention layers of the LLM, which may limit its direct applicability in certain settings where such access is unavailable, such as in some commercial API-driven models (e.g., GPT-4 OpenAI et al. (2023) and DeepSeek-R1 Guo et al. (2025)). Additionally, while our evaluations provide a solid foundation, further research is needed to explore the performance of Data Whisperer with larger LLMs and datasets. Expanding experiments to include larger and more complex models would offer valuable insights into the scalability of the method and its potential in extreme-scale scenarios. Addressing these areas in future work will help maximize the versatility and impact of Data Whisperer across a broader range of applications.
Appendix A More Experimental Details
A.1 Datasets
We use three real-world datasets for corresponding downstream tasks, namely, the BioInstruct dataset (Tran et al., 2024) for biomedical question answering, the DialogSum dataset (Chen et al., 2021) for dialogue summarization, and GSM8K (Cobbe et al., 2021) for mathematical reasoning. The BioInstruct dataset contains 25,005 instruction-input-output triplets spanning diverse biomedical scenarios, including clinical decision-making, biomedical question answering, and diagnostic interpretation. We split the dataset into training and testing sets with a 9:1 ratio. The DialogSum dataset is a comprehensive collection of 13,460 dialogues curated from open dialogue repositories, addressing a variety of everyday scenarios. Each dialogue is paired with a carefully constructed reference summary, enabling robust evaluation of summarization models. The GSM8K dataset offers 8.5K linguistically diverse, high-quality grade school math word problems requiring multi-step arithmetic reasoning. This dataset is designed to facilitate research on multi-step reasoning in question answering tasks for fundamental mathematical problem solving.
A.2 In-Context Learning Prompt Template
Figure 5: ICL prompt of GSM8k dataset.
Figure 6: ICL prompt of DialogSum dataset.
Figure 7: ICL prompt of BioInstruct dataset.
We formalize the design of In-Context Learning templates for three datasets: BioInstruct, DialogSum, and GSM8K. Each template is structured to provide explicit guidance through task-specific instructions, demonstrations, and output constraints.
Template Design. Our framework employs a unified template structure comprising three core components: (i) Task Instruction. A declarative statement that explicitly defines the task objective (e.g., "Generate a concise summary of the dialogue") alongside specific constraints (e.g., brevity, precision). (ii) In-Context Demonstrations. A set of input-output pairs that exemplify valid task execution, ensuring alignment with the task’s requirements and constraints. (iii) Output Constraints. Rigorous formatting rules that enforce compliance with task-specific syntax, such as step-by-step derivations for GSM8K or speaker-aware summarization for DialogSum. This modular architecture ensures consistency and adaptability across diverse tasks while maintaining strict adherence to domain-specific guidelines.
Task-Specific Prompt Design. We present a systematic approach to designing In-Context Learning prompts tailored to diverse datasets. For BioInstruct, we prioritize biomedical accuracy by implementing hierarchical guidelines that emphasize logical ordering and scope limitation, while ensuring responses are strictly evidence-based and devoid of speculative content. In DialogSum, we enforce structured dialogue processing through explicit speaker tagging (e.g., #Person1#), discourse compression techniques, and the generation of neutral third-person summaries. For GSM8K, we mandate a rigorous format that includes intermediate arithmetic steps and requires final answers to be boxed (e.g., #### <number>), explicitly prohibiting textual explanations to maintain precision and clarity. This tailored approach ensures that each dataset’s unique characteristics are effectively leveraged to optimize In-Context Learning performance. Structural schematics and formatting rules are visualized in Figure 5,  6 and  7.
A.3 Synthetic Data Generation
Figure 8: Data synthesis prompt of DialogSum dataset.
To evaluate the robustness of Data Whisperer on synthetic data, we constructed a synthetic variant of the DialogSum dataset using the Llama-3-8B-Instruct model. The generation pipeline comprised three key stages: prompt design, generation process, and quality control.
Prompt Design. For every 5 demonstration samples from the original DialogSum dataset, we crafted structured prompts to generate 1 synthetic dialogue-summary pair. These prompts enforced strict formatting rules, including speaker tagging and topic alignment, alongside explicit constraints to ensure the synthetic data closely mirrored the style and structure of the original dataset.
Generation Process. The model was instructed to produce dialogues with multiple turns, ensuring natural conversational flow and coherent narratives. Each output adhered to the JSON schema, with topics covering diverse everyday scenarios such as travel, shopping, and education.
Quality Control. A two-step filtering process was employed to ensure data quality. First, automated checks removed duplicates and syntax-invalid entries using regex-based validation. Second, three annotators independently reviewed the samples, flagging those with coherence issues, factual inconsistencies, or formatting violations. Disagreements were resolved through majority voting, and only samples free from hallucination or structural deviations were retained.
The final synthetic dataset maintained the original DialogSum distribution in dialogue length and topic diversity. Full prompt template is illustrated in Figure 8.
A.4 Hyper-Parameters of Different Methods
Data Selection. In our experiments, Data Whisperer was configured with task-specific batch sizes: 15 for the BioInstruct dataset, 5 for the DialogSum dataset, and 15 for the GSM8K dataset. By default, $n_{d}$ (number of demonstrations) was set to 10, and $n_{q}$ (number of queries) was set to 5. The generation temperature was fixed at 0 during the In-Context Learning process to ensure deterministic outputs. To ensure fairness, the batch size for all baseline methods was uniformly set to 16 across all tasks.
Model Fine-tuning. All models, whether trained on pruned or full datasets, shared identical hyperparameters. We employed parameter-efficient fine-tuning via LoRA Hu et al. (2021) for all tasks. The fine-tuning framework, based on Zheng et al. (2024) [^2] https://github.com/hiyouga/LLaMA-Factory, utilized a learning rate scheduler with linear warm-up and cosine decay. A consistent batch size of 8 was applied across all experiments. For Llama-3-8B-Instruct, the learning rate was fixed at $1\times 10^{-4}$ for all datasets. For Qwen-2.5-7B-Instruct and Mistral-Nemo-Instruct-2407, the learning rate was uniformly set to $1\times 10^{-5}$. All models were fine-tuned for 5 epochs, with each experiment repeated three times to ensure robustness and statistical reliability.
A.5 Pseudo Code of Data Whisperer
The detailed pseudo code of Data Whisperer is demonstrated in Algorithm 1.
Algorithm 1 Data Whisperer for Data Selection
Input:
$\mathcal{M}_{p}$ – Pre-trained language model
$\mathcal{D}$ – Full task dataset
$n_{d}$ – Number of samples for demonstration at a time
$n_{q}$ – Number of samples for query at a time
$I$ – Instruction for ICL generation
$f$ – Performance metric
$\rho$ – Budget ratio
Output: $\mathcal{D^{\prime}}$ – Selected coreset ($\mathcal{D^{\prime}}\subseteq\mathcal{D}$, $|\mathcal{D^{\prime}}|<|\mathcal{D}|$)
1:procedure DataWhisperer($\mathcal{M}_{p},\mathcal{D},n_{d},n_{q},I,f,\rho$)
2:     $\mathcal{S}\leftarrow\mathbf{0}$triangleright Initialize the score set
3:     while Not all samples in $\mathcal{D}$ have been processed do
4:         $\mathcal{D}_{d}\leftarrow\text{Randomly select }n_{d}\text{ samples from }\mathcal{D}$
5:         $\mathcal{D}_{q}\leftarrow\text{Randomly select }n_{q}\text{ samples from }\mathcal{D}$
6:         $C\leftarrow\{I,\mathcal{D}_{d}\}$triangleright Form input context
7:         $\hat{y}_{q}^{(1)},\dots,\hat{y}_{q}^{(n_{q})}\leftarrow\mathcal{M}_{p}(C,\mathcal{D}_{q})$triangleright Generate query predictions via ICL
8:         $s\leftarrow\frac{1}{n_{q}}\sum_{j=1}^{n_{q}}f(\hat{y}_{q}^{(j)},y_{q}^{(j)})$
9:         for $i=1\to n_{d}$ do
[…]
11:              $w_{i}\leftarrow\sum_{h}\mathbf{1}^{\top}A_{(x_{d}^{(i)},y_{d}^{(i)})}^{(h)}\mathbf{1}$triangleright Compute weights
12:         end for
13:         $w^{*}\leftarrow\text{Normalize}(w)$
14:         $s^{*}\leftarrow s\odot w^{*}$triangleright Element-wise multiplication with attention weights
15:         $\mathcal{S}[\mathcal{D}_{d}]\leftarrow\mathcal{S}[\mathcal{D}_{d}]+s^{*}$triangleright Accumulate scores for corresponding indices in $\mathcal{D}_{d}$
16:     end while
17:     $k\leftarrow\lfloor\rho\cdot|\mathcal{D}|\rfloor$
18:     $\mathcal{D}^{\prime}\leftarrow\text{Top-k}(\mathcal{D};\mathcal{S})$triangleright Select top $m$ samples based on scores
19:     return $\mathcal{D^{\prime}}$
20:end procedure
Table 6: Additional results on the Selection-to-Tuning ratio and performance scores for the Qwen and Mistral models are presented. The symbols $\uparrow$ and $\downarrow$ indicate improvements and degradations compared to random selection, respectively. “Speedup” refers to the acceleration achieved by Data Whisperer (w2s) over Nuggets. “w2s” denotes using the weaker model within the same family for data selection across each LLM.
Model Method GSM8K DialogSum BioInstruct
STR Performance STR Performance STR Performance
1% 5% 10% Avg. Score 1% 5% 10% Avg. Score 1% 5% 10% Avg. Score
Qwen-2.5-7B-Instruct Random - - - 60.29 - - - 34.66 - - - 35.19
GraNd 1.06 1.06 1.07 62.61$\uparrow$2.32 1.09 1.11 1.13 36.29$\uparrow$1.63 1.07 1.08 1.34 32.62$\downarrow$2.57
EL2N 1.06 1.06 1.07 60.72$\uparrow$0.43 1.10 1.13 1.15 36.97$\uparrow$2.31 1.08 1.08 1.32 34.64$\downarrow$0.55
CCS 1.04 1.04 1.05 61.34$\uparrow$1.05 1.01 1.04 1.06 35.30$\uparrow$0.64 1.02 1.03 1.25 31.87$\downarrow$3.32
Nuggets 1.10 1.11 1.11 62.74$\uparrow$2.45 3.66 3.69 3.71 36.63$\uparrow$1.97 2.19 2.20 2.41 34.95$\downarrow$0.24
STAFF 1.05 1.05 1.06 62.74$\uparrow$2.45 1.06 1.08 1.10 37.18$\uparrow$2.52 1.06 1.06 1.33 32.97$\downarrow$2.22
Data Whisperer 0.08 0.08 0.08 64.21$\uparrow$3.92 0.16 0.19 0.29 38.97$\uparrow$4.31 0.15 0.27 0.46 37.12$\uparrow$1.93
Data Whisperer (w2s) 0.05 0.05 0.06 63.67$\uparrow$3.38 0.10 0.13 0.23 38.37$\uparrow$3.71 3.71 0.20 0.37 35.98$\uparrow$0.79
Speedup 21.52$\times$ 21.12$\times$ 19.29$\times$ - 35.67$\times$ 27.48$\times$ 16.12$\times$ - 25.12$\times$ 10.94$\times$ 6.48$\times$ -
Mistral-Nemo-2407-12B Random - - - 50.36 - - - 32.61 - - - 25.00
GraNd 1.08 1.09 1.24 51.65$\uparrow$1.29 1.05 1.07 1.09 32.72$\uparrow$0.11 1.13 1.23 1.40 24.35$\downarrow$0.65
EL2N 1.11 1.13 1.36 51.74$\uparrow$1.38 1.06 1.09 1.11 30.75$\downarrow$1.86 1.14 1.23 1.39 23.76$\downarrow$1.24
CCS 1.04 1.01 1.22 50.71$\uparrow$0.35 1.01 1.04 1.05 31.75$\downarrow$0.86 1.02 1.11 1.25 26.37$\uparrow$1.37
Nuggets 2.34 2.35 2.50 53.06$\uparrow$2.70 2.22 2.24 2.26 32.53$\downarrow$0.08 5.15 5.23 5.36 28.12$\uparrow$3.12
STAFF 1.06 1.08 1.22 54.06$\uparrow$3.70 1.04 1.06 1.07 32.71$\uparrow$0.10 1.10 1.20 1.39 25.03$\uparrow$0.03
Data Whisperer 0.24 0.26 0.30 57.62$\uparrow$7.26 0.08 0.12 0.25 35.45$\uparrow$2.84 0.24 0.39 0.65 29.96$\uparrow$4.96
Data Whisperer (w2s) 0.11 0.12 0.12 57.21$\uparrow$6.85 0.07 0.11 0.24 35.07$\uparrow$2.46 0.18 0.33 0.60 29.66$\uparrow$4.66
Speedup 20.52$\times$ 20.45$\times$ 20.89$\times$ - 31.75$\times$ 20.69$\times$ 9.53$\times$ - 29.18$\times$ 15.61$\times$ 8.95$\times$ -
Appendix B Extensive Evaluation Results of STR
As defined in Eq. (1), the Selection-to-Tuning Ratio (STR) quantifies the efficiency of data selection methods by comparing the computational cost of selection to the time saved during fine-tuning. A method $\tau$ is efficient if $\text{STR}<1$, indicating that the time invested in selection is offset by faster fine-tuning. An STR below 1 ensures that the selection process positively contributes to overall pipeline efficiency, making it a practical and scalable choice for optimizing model training workflows. More evaluation results of STR are demonstrated in Table 6.